Startup Night 2 - 20140831 UT

Observers: JHill (LBTO), OKuhn, BRothberg (Tucson)
Telescope Operator: SAllanson (LBTO)
AO Support: JCGuerra, MLefevbre (LBTO)
Telescope Support: KNewton
Software Support: KSummers, DSummers (Tucson)
Instrument Support: JPower (LBTO)
Engineering Support: DAshby


PMH-9 power switch is OK.

LBC-Red shutter still has problems.

SX AdSec still will not turn on, but TSS module runs OK in the lab.

LBC guiding has been fixed. LBC-Blue filter wheel is working normally.

ARGOS electronics box has a minor glycol leak.


There were 6 power switches to generator and back before 2PM. MGIO worked out their procedural issues, and automatic switching is expected to be reliable.

Mike L and JC swapped in a spare TSS power module into the SX AdSec electronics, and there was no change. They tested the removed unit on the DX secondary in the lab, and found it to work fine (Both units work in the lab, neither one works on the telescope. The voltages in the power connector measure the same in both places.)

Kellee has fixed the LBC software bug that was stopping us from guiding. It was crashing in the rotational guiding routine that we never use.


Sky is photometric, wind is light. We are closed working on 4 technical problems.

a) Jenny and JRiedl (remote) are debugging the LBC-Red shutter. They are going to have to dismount the cryostat tomorrow.

b) Mike L and JC are continuing to debug the SX AdSec power problem.

c) Kellee is working on LBC software related to the UI.

d) Mike L checked the glycol leak in the ARGOS birdhouse. Inside the birdhouse was dry, so it seems to be a pinhole in the hose.


04:20 Open

04:53 N6946focus.ob - first one crashed.


05:05 Barry's dithered version on N6946 to test both dithering and guiding. Guiding works OK. Seeing is about 1 arcsec. The first dither worked

Changed executable link for PCS from PCS.2014C to PCS.tpk2.3.

06:03 stop this OB after 1 hour.

Stop and restart PCS to get the new pointing kernel.

06:05 Barry's dithered version on N6946 to test the new kernel (we did not refocus).

Stopped to Put the SX AdSec cabinet back together.

06:44 Preset to NCG7727


06:52 NGC7727_dither - seeing isn't too good down here. There was a huge (2 arcsec) jump in guiding/tracking just before 7UT. It appears on lbcb.20140831.065644.fits. We made Issue 5283 since we don't understand the cause.

JC and ML depart.

2014-08-31T06:58:50.485288+00:00 tcs2 LBT_PCS: Side: left Centroid Xrad: 0.000000041963709 Yrad: 0.000000321012807 Xarcs: 0.008655636249858 Yarcs: 0.066213644370860
2014-08-31T06:59:04.241859+00:00 tcs2 LBT_PCS: Side: left Centroid Xrad: 0.000000205530822 Yrad: 0.000000446206644 Xarcs: 0.042393775243173 Yarcs: 0.092036726952610
2014-08-31T06:59:43.673602+00:00 tcs2 LBT_PCS: Side: left Centroid Xrad: 0.000000801862426 Yrad: 0.000009925775330 Xarcs: 0.165395998026344 Yarcs: 2.047338125273574
2014-08-31T06:59:57.303961+00:00 tcs2 LBT_PCS: Side: left Centroid Xrad: -0.000000059358172 Yrad: -0.000008466870842 Xarcs: -0.012243501864550 Yarcs: -1.746417473818300

The lbc guider images show teh progression of this jump (lbcbtec.20140831.xxxxxx_1.fits):
065852_1.fits (looks good, normal)
065906 (star blew up to donut)
065918 (donut)
065931 (star coming back to focus)
065945 (pretty much OK by now)

07:30 DOFPIA - seeing is much better now.

07:35 starting the second OB on NGC7727 with modified exposure times and dither positions.


08:14 starting a third OB on NGC7727. images are 081453...083701

08:44 Slew to Landolt standard star field, SA92 (SA92_253), and observe at all blue filters.

09:00 Slew to USNO star near next field to collimate and correct pointing (using ptautoadjust)

09:18 Start observations of this field, Arp295.
  • from the start, but images were elongated and there is a gradient in FWHM, with FWHM increasing from top to bottom.

09:45 OB ended. Run dofpia again (dofpia, /x2, /blueonly).
  • at 09:59, dofpia hung... weird, I've never seen this happen. It reported "waiting for OB to finish..." but, while the OB did finish, and the last file was written to newdata (095901), displayed in RB_Science and is of the correct size, dofpia was paralyzed. Killed the idl process and restarted IDL and dofpia, and it appears to be working again.

10:10 Starting Science OB (3 dithers, 3 filters at each dither) PA=45 deg. Arp295.
  • At 0,0 dither position, the tech chip images show a star, but the thumbnail images at some point (during the U exposure) did not show a star - Star does not look that faint, but may have been below threshold for guiding?
  • Images still appear elongated from the start, despite repeating dofpia.

10:52 Repeating observation of SA92 field, now a little lower than earlier.

11:16 New field - N672.
  • dofpia, /blueonly
  • run OB - seeing is ~1".
  • completed in twilight (just before 5am)

11:58 Slew to SA92 again, now much lower than earlier, with airmass 1.457.

12:30 Tried some test flats, but it was already too bright. Steve is closing.

13:00 MODS calibrations to compare with pre-summer-shutdown. Dome dark, telescope at zenith.
  • bias8K
  • grpixflats
  • grlamps
  • grflats at 1.2" (since instconfig dual grating already done, can run only 1.2" block: execMODS -b LS60x1.2

-- %USERSIG{JohnHill - 2014-08-31}%
Topic revision: r7 - 31 Aug 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback